# Parametric Scaling Law of Tuning Bias in Conformal Prediction

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
In this work, we reveal a previously unrecognized phenomenon: the tuning bias - the coverage gap introduced by using the same dataset for tuning and calibration, is negligible for simple parameter tuning in conformal prediction methods. Empirically, we find that most methods maintain their coverage rates, except for vector scaling (Guo et al., 2017) and the fine-tuning version of ConfTr (Stutz et al., 2022). Furthermore, we observe the parametric scaling law of the tuning bias: this bias increases with parameter space complexity and decreases with calibration set size. An intuitive explanation for this phenomenon is that models with more complex parameter space require more data to tune for the optimal parameters.
@@ 2 Background
For the test step with a given feature of test sample $\boldsymbol{x}$, the non-conformity score is computed for each label $y\in\mathcal{Y}$. The prediction set $\widehat{\mathcal{C}}(\boldsymbol{x})$ is constructed by including all labels whose non-conformity scores are below the threshold $\hat{t}$:
$\widehat{\mathcal{C}}(\boldsymbol{x}):=\widehat{\mathcal{C}}_{\hat{t}}(\boldsymbol{x})=\left\{y\in\mathcal{Y}:S(\boldsymbol{x},y)\leq\hat{t}\right\}.$ (3)
By this construction, the conformal prediction set $\widehat{\mathcal{C}}(\boldsymbol{x})$ in Eq. (3) has a finite-sample coverage guarantee with the assumption of exchangeability:
@@
$1-\alpha\leq\mathbb{P}\left(y\in\widehat{\mathcal{C}}(\boldsymbol{x})\right)\leq 1-\alpha+\frac{1}{n+1}.$
@@
$\mathop{CovGap}(\mathcal{C}):=|(1-\alpha)-\mathbb{P}(y\in\mathcal{C}(\boldsymbol{x}))|.$
@@
For the conformal prediction with parameter tuning, the tuning bias for parameter tuning on the same set as the calibration set is defined as the additional coverage gap caused by the practice, i.e.,
$\mathop{TuningBias}(\widehat{\mathcal{C}})=\mathop{CovGap}(\widehat{\mathcal{C}})-\mathop{CovGap}(\widehat{\mathcal{C}}_{\text{hold-out}}),$
where $\widehat{\mathcal{C}}$ is the conformal predictor with parameter tuned on the same set as the calibration set, and $\widehat{\mathcal{C}}_{\text{hold-out}}$ is the conformal predictor with parameter tuning using a hold-out set separated from the set used for conformal prediction.
@@ 3 Empirical Study
Figure 1: Tuning biases of various methods in conformal prediction, using ResNet-18 on CIFAR-100 at $\alpha=0.1$. The APS is used except for RAPS, SAPS and SA, with a calibration set size of 1000. The ‘hold-out’ and ‘same’ denote that we use separate/same datasets for parameter tuning and calibration, respectively. Tuning bias is the difference between the coverage gaps of models tuned in the setting of hold-out and same. The tuning biases are not large, except for VS and ConfTr (ft.).
@@ 3.1 Tuning Bias of Current Methods
Tuning bias is negligible for simple parameter tuning.
From Figure 1, the tuning biases are not significant for most methods, except for vector scaling and the fine-tuning version of ConfTr. The tuning bias is not significant for simple parameter tuning as a result of the small size of the parameter space. The large tuning bias observed in vector scaling and the fine-tuning version of ConfTr may be due to the large number of parameters requiring tuning. This tuning could potentially result in overfitting, where the transformed score distribution is overly adapted to the hold-out set, significantly reducing its performance on the test sample. As the size of the calibration set increases, the tuning bias converges to zero because more data helps reduce the generalization bias. Following this, we further explore influences on the tuning bias from the perspectives of parameter space and calibration set.
@@ 3.2 Parametric Scaling Laws of Tuning Bias
We explore the factors influencing the tuning biases. By the convention to control the number of unfrozen parameters, we focus on confidence calibration using vector scaling. We could control the number of unfrozen parameters by freezing a portion of the parameters in vector scaling. We freeze a portion of the parameters in vector scaling. In detail, we do not let the proportion of the parameters be optimized randomly. The ratio of frozen parameters ranges from $0\%$ to $40\%$ by $10\%$ increments. The detail of vector scaling is provided in Section A.2. We use separate or the same datasets for parameter tuning and calibration, respectively, with APS. All methods of conformal prediction are under the same error rate $\alpha=0.1$. To ensure the reliability of our results, we repeated each experiment 30 times. And we reported both the average result and its corresponding standard deviation, respectively.
@@ 3.3 Intuitive Explanation
$\hat{t}$ $=\arg\min_{t}\sum_{i\in\mathcal{I}_{\text{cal}}}|\hat{\mathcal{C}}_{t}(\boldsymbol{x}_{i})|/n$
@@ 4.1 Problem Formulation
The conformal prediction set with parameter tuning is:
$\widehat{\mathcal{C}}(\boldsymbol{x}):=\widehat{\mathcal{C}}^{\hat{\lambda}}_{\hat{t}^{\hat{\lambda}}_{\mathcal{D}_{\text{cal}}}}(\boldsymbol{x}).$ (5)
@@
$\mathop{CovGap}(\widehat{\mathcal{C}},\alpha)\leq\mathbb{E}\mathfrak{R}_{\Lambda}+\varepsilon_{\alpha,n},$
@@
The proof of Theorem 4.1 is provided in Appendix D. The upper bound of the coverage gap could be decomposed into two components: $\mathbb{E}\mathfrak{R}_{\Lambda}$ and $\varepsilon_{\alpha,n}$. The first term, $\mathbb{E}\mathfrak{R}{\Lambda}$, introduces an additional bound for the coverage gap, which is not present in the standard conformal prediction framework, containing only the second term, $\varepsilon_{\alpha,n}$. As the definition of tuning bias in Definition 2.2, we could conclude that
$\mathop{TuningBias}(\widehat{\mathcal{C}})\leq\mathbb{E}\mathfrak{R}_{\Lambda}.$
@@
=$\mathcal{H}=\left\{\mathds{1}\{S^{\lambda}(\boldsymbol{x},y)\leq t\},\lambda\in\Lambda,t\in\mathbb{R}\right\}.$
The bound $\mathbb{E}\mathfrak{R}_{\Lambda}$ is the empirical process on the class $\mathcal{H}$:
$\mathbb{E}\mathfrak{R}_{\Lambda}=\mathbb{E}\sup_{g\in\mathcal{H}}\Big{|}\frac{1}{n}\sum_{i\in[n]}g(\boldsymbol{x}_{i},y_{i})-\mathbb{E}[g(\boldsymbol{x}_{\text{test}},y_{\text{test}})\mid\mathcal{D}_{\text{cal}}]\Big{|}.$
@@ 4.2 Tuning Bias in Finite Parameter Space
$\mathop{TuningBias}(\widehat{\mathcal{C}})\leq\sqrt{\frac{\log(2|\Lambda|)}{2n}}+\frac{1}{\sqrt{2n}\sqrt{\log(2|\Lambda|)}}.$
The proof of Proposition 4.2 is provided in Appendix E. The result in Proposition 4.2 is a general result that applies to any finite parameter space $\Lambda$. The examples include parameter tuning in some specific score functions (Angelopoulos et al., 2021; Huang et al., 2024), score selection (Yang & Kuchibhotla, 2024) and aggregation (Luo & Zhou, 2024) with finite candidates, and early stopping in conformal prediction (Liang et al., 2023) and so on.
@@
$\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{RAPS}})\leq\sqrt{\frac{\log(2M\bar{K})}{2n}}+\frac{1}{\sqrt{2n}\sqrt{\log(2M\bar{K})}}.$
The proof of Corollary 4.3 is provided in Appendix F. Although the RAPS method (Angelopoulos et al., 2021) recommends using an additional hold-out set for hyperparameter selection, our findings allow for direct hyperparameter selection without the need for an extra hold-out set. As for SAPS (Huang et al., 2024), the parameter space is also finite, and we could also apply the same analysis to SAPS.
@@
$\hat{m}=\arg\min_{m\in\{1,\ldots,M\}}\sum_{i\in\mathcal{I}_{\text{cal}}}|\mathcal{C}_{m}(\boldsymbol{x}_{i})|/|\mathcal{I}_{\text{cal}}|.$
@@
$\mathop{TuningBias}(\widehat{\mathcal{C}})\leq\sqrt{\frac{\log(2M)}{2n}}+\frac{1}{\sqrt{2n}\sqrt{\log(2M)}}.$
The proof of Corollary 4.4 is provided in Appendix F. For the score aggregation (Luo & Zhou, 2024), we analyze it in Appendix K. Next, we extend our analysis to an infinite parameter space and apply it to conformal prediction with confidence calibration, where the parameter space is infinite.
@@ 4.3 Tuning Bias in Infinite Parameter Space
Although the concept may seem hard to understand, it can be viewed as the maximum number of points that the class of functions can shatter intuitively. With a more complex class of functions, it is more likely to shatter a larger number of points. It is a measure of the complexity of the class of functions, $\mathcal{F}$. The intuition is that the larger the complexity of the class of functions, the larger the bound of the empirical process is. By the definition of $\mathcal{H}$ and Lemma C.6 in Appendix C, we develop a general bound of the tuning bias:
@@
$\mathop{TuningBias}(\widehat{\mathcal{C}})\leq C\sqrt{\frac{d+1}{n}}$
@@
The proof of Proposition 4.6 is provided in Appendix G. The bound in Proposition 4.6 is a general result that applies to any infinite parameter space $\Lambda\subset\mathbb{R}^{d}$. It shows that the tuning bias increases with the dimension of parameter space $d$ and decreases with the size of the hold-out set $n$. Next, we will present an application for confidence calibration.
@@
For a classification problem with $K$ classes, the calibration set size $n$ and the same pre-trained model, the tuning bias of temperature scaling is smaller than that of vector scaling: $\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{TS}})\leq\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{VS}})$, where $\widehat{\mathcal{C}}_{\text{TS}}$ and $\widehat{\mathcal{C}}_{\text{VS}}$ are the prediction sets of temperature scaling and vector scaling, respectively.
The proof of Corollary 4.7 is provided in Appendix H. The result shows that the tuning bias of temperature scaling is smaller than that of vector scaling. We provide a result for the binary classification case in Section 5, where the tuning bias of temperature scaling is zero.
@@ 5 Discussion of Potential Solutions
Methods CIFAR-100 ImageNet
Temperature scaling 0.14% 0.04%
Vector scaling 1.13% 6.63%
ConfTr (ft.) w/  OP 0.04% 0.02%
ConfTr (ft.) w/o OP 5.5% 21.3%
@@
A straightforward way to constrain parameter complexity is to reduce the number of parameters in model tuning. Here, we provide an example of parameter reduction through weight sharing. In particular, we show the benefits of weight sharing by comparing the tuning biases of temperature scaling (TS) and vector scaling (VS), where the former can be viewed as a special case of vector scaling with a shared parameter $|\Lambda_{\mathrm{TS}}|<|\Lambda_{\mathrm{VS}}|$. The empirical results in Table 1 show that TS with fewer parameters achieves much smaller tuning biases than VS. Theoretically, we provide a theoretical result for the comparison in the case of binary classification:
@@
For a binary classification problem, we have $\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{TS}})=0\leq\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{VS}})$, where $\widehat{\mathcal{C}}_{\text{TS}}$ and $\widehat{\mathcal{C}}_{\text{VS}}$ are the prediction sets of models tuned by temperature scaling and vector scaling, respectively.
We provide the proof of the above proposition in Appendix I. From the example, we demonstrate that designing tuning algorithms with few parameters will be beneficial in preventing the tuning bias. However, reducing the parameter number may severely degrade the performance of parameter tuning, making it inapplicable in some scenarios.
@@
For a classification problem with $K$ classes, and the calibration set size $n$, we have the tuning bias of $\widehat{\mathcal{C}}_{\text{op}}$ is less than the one of $\widehat{\mathcal{C}}$: $\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{op}})\leq\mathop{TuningBias}(\widehat{\mathcal{C}})$, where $\widehat{\mathcal{C}}$ and $\widehat{\mathcal{C}}_{\text{op}}$ represent the prediction sets of models tuned by ConfTr, with $\widehat{\mathcal{C}}_{\text{op}}$ under the order-preserving constraint and $\widehat{\mathcal{C}}$ without it.
The proof of Proposition 5.2 is provided in Appendix J. We also verify the argument by the empirical results shown in Table 1. The tuning bias of ConfTr with order-preserving constraint is much smaller than that without it. This approach is advantageous as it can generalize to complex paradigms of parameter tuning (e.g., fine-tuning the final layer). It is worth noting that designing tailored regularization techniques for specific tuning methods in conformal prediction may be a promising direction for future research.

### v2

@@ Abstract
Machine Learning, Uncertainty Quantification, Conformal Prediction, Parameter Tuning, Coverage Gap, Tuning Bias
@@ 1 Introduction
In this work, we reveal a previously unrecognized phenomenon: the tuning bias - the coverage gap introduced by using the same dataset for tuning and calibration, is negligible for simple parameter tuning in conformal prediction methods. Empirically, we find that most methods maintain their coverage rates, except for vector scaling (Guo et al., 2017) and the fine-tuning version of ConfTr (Stutz et al., 2022). Furthermore, we observe the parametric scaling law of the tuning bias: this bias increases with parameter space complexity and decreases with calibration set size. An intuitive explanation for this phenomenon is that models with more complex parameter space require more data to tune for the optimal parameters. [^1] Code: https://github.com/ml-stat-Sustech/Parametric-Scaling-Law-CP-Tuning.
@@ 2 Background
For the test step with a given feature of test sample $\boldsymbol{x}$, the non-conformity score is computed for each label $y\in\mathcal{Y}$. The prediction set ${\mathcal{C}}(\boldsymbol{x})$ is constructed by including all labels whose non-conformity scores are below the threshold $\hat{t}$:
${\mathcal{C}}(\boldsymbol{x}):=\widehat{\mathcal{C}}_{\hat{t}}(\boldsymbol{x})=\left\{y\in\mathcal{Y}:S(\boldsymbol{x},y)\leq\hat{t}\right\},$ (3)
where the notation $\widehat{\cdot}$ is used to emphasize its dependence on the calibration set $\mathcal{D}_{\text{cal}}$. Here we consider a little weaker version of independence:
Definition 2.1 (Exchangeability).
The samples from $\mathcal{D}$ are exchangeable if and only if the joint distribution of these samples is invariant under any permutation of the indices of the samples in $\mathcal{D}$.
With the assumption of exchangeability, the conformal prediction set ${\mathcal{C}}(\boldsymbol{x})$ in Eq. (3) has a finite-sample coverage guarantee:
@@
$1-\alpha\leq\mathbb{P}\left(y\in{\mathcal{C}}(\boldsymbol{x})\right)\leq 1-\alpha+\frac{1}{n+1}.$
@@
$\mathop{\text{CovGap}}(\mathcal{C}):=|(1-\alpha)-\mathbb{P}(y\in\mathcal{C}(\boldsymbol{x}))|.$
@@
For the conformal prediction with parameter tuning, $\mathcal{C}$, the tuning bias for parameter tuning on the same set as the calibration set is defined as the additional coverage gap caused by the practice, i.e.,
$\mathop{\text{TuningBias}}(\mathcal{C})=\mathop{\text{CovGap}}(\mathcal{C})-\mathop{\text{CovGap}}(\mathcal{C}_{\text{hold-out}}),$
where $\mathcal{C}$ is the conformal predictor with parameter tuned on the same set as the calibration set, and $\mathcal{C}_{\text{hold-out}}$ is the conformal predictor with parameter tuning using a hold-out set separated from the set used for conformal prediction.
Remark 2.4.
Here, we clarify that understanding the tuning bias is crucial in conformal prediction practice:
  • For data-scarce scenarios, splitting the calibration dataset is impractical in data-scarce scenarios like rare diseases, natural disaster prediction, and privacy-constrained personal data. With limited data, using separated datasets will reduce points for parameter tuning and conformal calibration, compromising the approach’s effectiveness and stability. Thus, it’s valuable to assess when splitting is needed, or data reuse is permissible rather than sticking to traditional practices.
  • For simple implementation, even with sufficient data, maintaining separate sets can increase the pipeline complexity. Understanding when this separation is unnecessary—such as when tuning bias is negligible—enables simpler, more streamlined workflows while preserving coverage guarantees, offering practical relevance.
  • For foundational understanding, exploring the tuning bias can provide an in-depth understanding of the exchangeability assumption in conformal prediction. In particular, the insight in this work may inspire future works in non-exchangeable conformal prediction.
It could also answer the concern, “why not split the dataset”, for the same reasons. Indeed, it is easy to split the validation set when the provided data is sufficient. However, it can be particularly important to consider data reusing in data-scarce scenarios, like rare diseases, natural disaster prediction, and privacy-constrained personal data. With limited data, separating the dataset will further exacerbate the data scarcity problem, compromising the effectiveness of both conformal calibration and parameter tuning. In addition, exploring the tuning bias could provide an in-depth understanding of the exchangeability assumption, which may inspire future works in non-exchangeable CP.
@@ 3 Empirical Study
Figure 1: Tuning biases of various methods in conformal prediction, using ResNet-18 on CIFAR-100 at $\alpha=0.1$. The APS is used except for RAPS, SAPS and SA, with a calibration set size of 1000. The “hold-out” and “same” denote that we use separate/same datasets for parameter tuning and calibration, respectively. Tuning bias is the difference between the coverage gaps of models tuned in the setting of hold-out and same.
@@ 3.1 Tuning Bias of Current Methods
Tuning bias is not always negligible for the parameter tuning in conformal prediction.
In Figure 1, we present the tuning biases caused by various tuning methods using ResNet-18 on CIFAR-100. The results indicate that most methods do not introduce significant tuning biases, achieving similar coverage gaps regardless of whether the same dataset or separate datasets are used for calibration and tuning. For example, the coverage gaps of RAPS using the same and separate datasets are almost identical, as shown in Figure 1. However, in vector scaling and ConfTr (ft.), using the same datasets could result in much larger coverage gaps than using a hold-out dataset for tuning. We conjecture that these differences may be due to the large number of parameters tuned in these methods. In what follows, we provide in-depth analyses to explore the factors influencing tuning biases in these methods.
@@ 3.2 Parametric Scaling Laws of Tuning Bias
We conduct an additional analysis from two key perspectives: (1) the number of parameters and (2) the size of the calibration set. The experiments provide deeper insights into how tuning methods influence bias formation. By the convention to control the number of unfrozen parameters, we focus on confidence calibration using vector scaling. We could control the number of unfrozen parameters by freezing a portion of the parameters in vector scaling. We freeze a portion of the parameters in vector scaling. In detail, we do not let the proportion of the parameters be optimized randomly. The ratio of frozen parameters ranges from $0\%$ to $40\%$ by $10\%$ increments. The detail of vector scaling is provided in Section A.2. We use separate or the same datasets for parameter tuning and calibration, respectively, with APS. All methods of conformal prediction are under the same error rate $\alpha=0.1$. To ensure the reliability of our results, we repeat each experiment 30 times and report both the average result and its corresponding standard deviation.
@@ 3.3 Intuitive Explanation
$\hat{t}$ $=\arg\min_{t}\sum_{i\in\mathcal{I}_{\text{cal}}}|\widehat{\mathcal{C}}_{t}(\boldsymbol{x}_{i})|/n$
@@ 4.1 Problem Formulation
The conformal prediction set with parameter tuning is reformulated as:
${\mathcal{C}}(\boldsymbol{x}):=\widehat{\mathcal{C}}^{\hat{\lambda}}_{\hat{t}^{\hat{\lambda}}_{\mathcal{D}_{\text{cal}}}}(\boldsymbol{x}).$ (5)
@@
$\mathop{\text{CovGap}}({\mathcal{C}})\leq\mathbb{E}\mathfrak{R}_{\Lambda}+\varepsilon_{\alpha,n},$
@@
The proof of Theorem 4.1 is provided in Appendix E. The upper bound of the coverage gap could be decomposed into two components: $\mathbb{E}\mathfrak{R}_{\Lambda}$ and $\varepsilon_{\alpha,n}$. The first term, $\mathbb{E}\mathfrak{R}{\Lambda}$, introduces an additional bound for the coverage gap, which is not present in the standard conformal prediction framework, containing only the second term, $\varepsilon_{\alpha,n}$. As the definition of tuning bias in Definition 2.3, we could conclude that
$\mathop{\text{TuningBias}}({\mathcal{C}})\leq\mathbb{E}\mathfrak{R}_{\Lambda}.$
@@
$\mathcal{H}_{\Lambda}=\left\{\mathds{1}\{S^{\lambda}(\boldsymbol{x},y)\leq t\},\lambda\in\Lambda,t\in\mathbb{R}\right\}.$
The bound $\mathbb{E}\mathfrak{R}_{\Lambda}$ is the empirical process on the class $\mathcal{H}_{\Lambda}$:
$\mathbb{E}\mathfrak{R}_{\Lambda}=\mathbb{E}\sup_{g\in\mathcal{H}_{\Lambda}}\Big{|}\frac{1}{n}\sum_{i\in[n]}g(\boldsymbol{x}_{i},y_{i})-\mathbb{E}[g(\boldsymbol{x}_{\text{test}},y_{\text{test}})\mid\mathcal{D}_{\text{cal}}]\Big{|}.$
@@ 4.2 Tuning Bias in Finite Parameter Space
$\mathop{\text{TuningBias}}({\mathcal{C}})\leq\sqrt{\frac{\log(2|\Lambda|)}{2n}}+\frac{1}{\sqrt{2n}\sqrt{\log(2|\Lambda|)}}.$
The proof of Proposition 4.2 is provided in Appendix F. The result in Proposition 4.2 is a general result that applies to any finite parameter space $\Lambda$. The examples include parameter tuning in some specific score functions (Angelopoulos et al., 2021; Huang et al., 2024), score selection (Yang & Kuchibhotla, 2024) and aggregation (Luo & Zhou, 2024) with finite candidates, and early stopping in conformal prediction (Liang et al., 2023) and so on.
@@
$\mathop{\text{TuningBias}}({\mathcal{C}}_{\text{RAPS}})\leq\sqrt{\frac{\log(2M\bar{K})}{2n}}+\frac{1}{\sqrt{2n}\sqrt{\log(2M\bar{K})}}.$
The proof of Corollary 4.3 is provided in Appendix G. Although the RAPS method (Angelopoulos et al., 2021) recommends using an additional hold-out set for hyperparameter selection, our findings allow for direct hyperparameter selection without the need for an extra hold-out set. As for SAPS (Huang et al., 2024), the parameter space is also finite, and we could also apply the same analysis to SAPS.
@@
$\hat{m}=\underset{m\in\{1,\ldots,M\}}{\operatorname{arg\,min\,}}\sum_{i\in\mathcal{I}_{\text{cal}}}|\mathcal{C}_{m}(\boldsymbol{x}_{i})|/|\mathcal{I}_{\text{cal}}|.$
@@
$\mathop{\text{TuningBias}}({\mathcal{C}})\leq\sqrt{\frac{\log(2M)}{2n}}+\frac{1}{\sqrt{2n}\sqrt{\log(2M)}}.$
The proof of Corollary 4.4 is provided in Appendix G. For the score aggregation (Luo & Zhou, 2024), we analyze it in Appendix L. Next, we extend our analysis to an infinite parameter space and apply it to conformal prediction with confidence calibration, where the parameter space is infinite.
@@ 4.3 Tuning Bias in Infinite Parameter Space
Although the concept may seem hard to understand, it can be viewed as the maximum number of points that the class of functions can shatter intuitively. With a more complex class of functions, it is more likely to shatter a larger number of points. It is a measure of the complexity of the class of functions, $\mathcal{F}$. The intuition is that the larger the complexity of the class of functions, the larger the bound of the empirical process is. By the definition of $\mathcal{H}$ and Lemma D.6 in Appendix D, we develop a general bound of the tuning bias:
@@
$\mathop{\text{TuningBias}}(\widehat{\mathcal{C}})\leq C\sqrt{\frac{d+1}{n}}$
@@
The proof of Proposition 4.6 is provided in Appendix H. The bound in Proposition 4.6 is a general result that applies to any infinite parameter space $\Lambda\subset\mathbb{R}^{d}$. It shows that the tuning bias increases with the dimension of parameter space $d$ and decreases with the size of the hold-out set $n$. Next, we will present an application for confidence calibration.
@@
For a classification problem with $K$ classes, the calibration set size $n$ and the same pre-trained model, the tuning bias of temperature scaling is smaller than that of vector scaling: $\mathop{\text{TuningBias}}({\mathcal{C}}_{\text{TS}})\leq\mathop{\text{TuningBias}}({\mathcal{C}}_{\text{VS}})$, where ${\mathcal{C}}_{\text{TS}}$ and ${\mathcal{C}}_{\text{VS}}$ are the prediction sets of temperature scaling and vector scaling, respectively.
The proof of Corollary 4.7 is provided in Appendix I. The result shows that the tuning bias of temperature scaling is smaller than that of vector scaling. We provide a result for the binary classification case in Section 5, where the tuning bias of temperature scaling is zero.
@@ 5 Discussion of Potential Solutions
Methods CIFAR-100 (%) ImageNet (%)
Temperature scaling 0.14 $\pm$ 0.01 0.04 $\pm$ 0.01
Vector scaling 1.13 $\pm$ 0.02 6.63 $\pm$ 0.03
ConfTr (ft.) w/ OP 0.52 $\pm$ 0.37 0.40 $\pm$ 0.31
ConfTr (ft.) w/o OP 6.15 $\pm$ 0.86 21.68 $\pm$ 0.58
@@
A straightforward way to constrain parameter complexity is to reduce the number of parameters in model tuning. Here, we provide an example of parameter reduction through weight sharing. In particular, we show the benefits of weight sharing by comparing the tuning biases of temperature scaling (TS) and vector scaling (VS), where the former can be viewed as a special case of vector scaling with a shared parameter $|\Lambda_{\mathrm{TS}}|<|\Lambda_{\mathrm{VS}}|$. The empirical results in Table 1 show that TS with fewer parameters achieves much smaller tuning biases than VS, which is also supported by Figure 1 and Figure 4. Theoretically, we provide a theoretical result for the comparison in the case of binary classification:
@@
For a binary classification problem, we have $\mathop{\text{TuningBias}}({\mathcal{C}}_{\text{TS}})=0\leq\mathop{\text{TuningBias}}({\mathcal{C}}_{\text{VS}})$, where ${\mathcal{C}}_{\text{TS}}$ and ${\mathcal{C}}_{\text{VS}}$ are the prediction sets of models tuned by temperature scaling and vector scaling, respectively.
We provide the proof of the above proposition in Appendix J. From the example, we demonstrate that designing tuning algorithms with few parameters will be beneficial in preventing the tuning bias. Also, we analyze the relationship between the number of parameters and tuning bias by freezing different numbers of parameters within VS. The results in Figure 2 (a) show that increasing the number of parameters leads to higher tuning bias, supporting the claim. However, reducing the parameter number may severely degrade the performance of parameter tuning, making it inapplicable in some scenarios, which motivates us to apply regularization to the tuning algorithm.
@@
For a classification problem with $K$ classes, and the calibration set size $n$, we have the tuning bias of ${\mathcal{C}}_{\text{op}}$ is less than the one of ${\mathcal{C}}$: $\mathop{\text{TuningBias}}({\mathcal{C}}_{\text{op}})\leq\mathop{\text{TuningBias}}({\mathcal{C}})$, where ${\mathcal{C}}$ and ${\mathcal{C}}_{\text{op}}$ represent the prediction sets of models tuned by ConfTr, with ${\mathcal{C}}_{\text{op}}$ under the order-preserving constraint and ${\mathcal{C}}$ without it.
The proof of Proposition 5.2 is provided in Appendix K. And by Lemma K.1, the dimension of parameter space of order-preserving vector scaling is $2$, which is much smaller than that without constraint, $2K$. By Lemma K.3, the dimension of parameter space of ConfTr with order-preserving constraint is $K+2$, which is much smaller than that of matrix scaling with its dimension being $K^{2}+K$. We also verify the argument by the empirical results shown in Table 1. The tuning bias of ConfTr with order-preserving constraint is much smaller than that without it. This approach is advantageous as it can generalize to complex paradigms of parameter tuning (e.g., fine-tuning the final layer). It is worth noting that designing tailored regularization techniques for specific tuning methods in conformal prediction may be a promising direction for future research.